# Divide, Interact, Sample: The Two-System Paradigm

### v1

@@ A.4 Propagation of Chaos
$\begin{split}dX^{i}_{t}\ &=\ b(t,X^{i}_{t},\delta_{\mathbf{Y}_{t}})\,dt\ +\ \sigma(t,X^{i}_{t},\delta_{\mathbf{Y}_{t}})\,dW^{i}_{t}\\ dY^{i}_{t}\ &=\ b(t,Y^{i}_{t},\delta_{\mathbf{X}_{t}})\,dt\ +\ \sigma(t,Y^{i}_{t},\delta_{\mathbf{X}_{t}})\,dW^{i}_{t}\\ d\overline{X}^{i}_{t}\ &=\ b(t,\overline{X}^{i}_{t},\overline{\mu}_{t})\,dt\ +\ \sigma(t,X^{i}_{t},\overline{\mu}_{t})\,dW^{i}_{t},\end{split}$ (80)
for $i\in\{1,\ldots,N\}$, with identical initial conditions $X^{i}_{0}=Y^{i}_{0}=\overline{X}^{i}_{0}=X_{0}$, we have
$\lim_{N\to\infty}\sup_{1\leq i\leq N}\mathbb{E}\left(\sup_{0\leq t\leq T}\lvert X^{i}_{s}-\overline{X}^{i}_{s}\rvert^{2}+\sup_{0\leq t\leq T}\lvert Y^{i}_{s}-\overline{X}^{i}_{s}\rvert^{2}\right)\ =\ 0,$ (81)
@@
$\sup_{1\leq i\leq N}\left(\mathbb{E}\sup_{0\leq t\leq T}\lvert X^{i}_{t}-\overline{X}^{i}_{t}\rvert^{2}\ +\ \mathbb{E}\sup_{0\leq t\leq T}\lvert Y^{i}_{t}-\overline{X}^{i}_{t}\rvert^{2}\right)\ \leq\ \frac{c}{N},$ (82)
@@
$\leq\ c\int^{t}_{0}\mathbb{E}\lvert b(s,X^{i}_{s},\delta_{\mathbf{Y}_{s}})-b(s,\overline{X}^{i}_{s},\delta_{\overline{\mathbf{X}}_{s}})\rvert^{2}\,ds$ (85)
$\qquad\qquad+\ c\int^{t}_{0}\mathbb{E}\lvert b(s,\overline{X}^{i}_{s},\delta_{\overline{\mathbf{X}}_{s}})-b(s,\overline{X}^{i}_{s},\overline{\mu}_{s})\rvert^{2}\,ds$ (86)
$+\ c\int^{t}_{0}\mathbb{E}\lvert\sigma(s,X^{i}_{s},\delta_{\mathbf{Y}_{s}})-\sigma(s,\overline{X}^{i}_{s},\delta_{\overline{\mathbf{X}}_{s}})\rvert^{2}\,ds$ (87)
$\qquad\qquad+\ c\int^{t}_{0}\mathbb{E}\lvert\sigma(s,\overline{X}^{i}_{s},\delta_{\overline{\mathbf{X}}_{s}})-\sigma(s,\overline{X}^{i}_{s},\overline{\mu}_{s})\rvert^{2}\,ds,$ (88)
with an similar inequality holding for $\mathbb{E}\sup_{0\leq s\leq t}\lvert Y^{i}_{s}-\overline{X}^{i}_{s}\rvert^{2}$.
The Wasserstein distance between two Dirac measures is bounded by (Equation 1.24 [7])
$\mathcal{W}_{2}^{2}(\delta_{\mathbf{Y}_{s}},\delta_{\overline{\mathbf{X}}_{s}})\ \leq\ \frac{1}{N}\sum_{i=1}^{N}\lvert Y_{s}^{i}-\overline{X}_{s}^{i}\rvert$. (89)
By the exchangeability of the couples $(Y^{i},\overline{X}^{i})$, the Lipschitz assumption of $b$ and $\sigma$ bounds the first and third term of the above right-hand side by
$c\int^{t}_{0}\left(\mathbb{E}\sup_{0\leq r\leq s}\lvert X_{r}^{i}-\overline{X}^{i}_{r}\rvert^{2}\ +\ \mathbb{E}\sup_{0\leq r\leq s}\lvert Y^{i}_{r}-\overline{X}^{i}_{r}\rvert^{2}\right)\,ds.$ (90)
Then using Grönwall’s lemma yields and the Lipschitz assumption
$\mathbb{E}\sup_{0\leq t\leq T}\lvert X^{i}_{t}-\overline{X}^{i}_{t}\rvert^{2}\ +\ \mathbb{E}\sup_{0\leq t\leq T}\lvert Y^{i}_{t}-\overline{X}^{i}_{t}\rvert^{2}\$$\leq\ c\int^{T}_{0}\mathbb{E}\lvert b(s,\overline{X}^{i}_{s},\delta_{\overline{\mathbf{X}}_{s}})-b(s,\overline{X}^{i}_{s},\overline{\mu}_{s})\rvert^{2}\,ds$ (91)
@@
$\leq c\int^{T}_{0}\mathbb{E}\mathcal{W}_{2}^{2}(\delta_{\overline{\mathbf{X}}_{s}},\overline{\mu}_{s})\,ds.$ (93)
Since $\overline{X}_{s}^{i}$ are all independent random variables with common law $\overline{\mu}_{s}$, it follows by Lemma 2 and Lebesgue’s dominated convergence theorem that
$\lim_{N\to\infty}\sup_{1\leq i\leq N}\mathbb{E}\left(\sup_{0\leq t\leq T}\lvert X^{i}_{t}-\overline{X}^{i}_{t}\rvert^{2}+\sup_{0\leq t\leq T}\lvert Y^{i}_{t}-\overline{X}^{i}_{t}\rvert^{2}\right)\ =\ 0.$ (94)
@@
$\qquad\qquad\qquad\qquad\cdot\bigg{[}\int b(s,\overline{X}^{i}_{s},y)\overline{\mu}_{s}(dy)-b(s,\overline{X}^{i},\delta_{\overline{X}^{l}_{s}})\bigg{]}\bigg{)}.$ (97)
@@
$\sup_{1\leq i\leq N}\left(\mathbb{E}\sup_{0\leq t\leq T}\lvert X^{i}_{t}-\overline{X}^{i}_{t}\rvert^{2}\ +\ \mathbb{E}\sup_{0\leq t\leq T}\lvert Y^{i}_{t}-\overline{X}^{i}_{t}\rvert^{2}\right)\ \leq\ \frac{c}{N}.$ (98)
@@ A.5 Adaptive MALA
The Metropolis-Hastings acceptance ratio is chosen as

### v2

@@ A.4 Propagation of Chaos
$\begin{split}dX^{i}_{t}\ &=\ b(t,X^{i}_{t},\delta_{\mathbf{Y}_{t}})\,dt\ +\ \sigma(t,X^{i}_{t},\delta_{\mathbf{Y}_{t}})\,dW^{X,i}_{t}\\ dY^{i}_{t}\ &=\ b(t,Y^{j}_{t},\delta_{\mathbf{X}_{t}})\,dt\ +\ \sigma(t,Y^{i}_{t},\delta_{\mathbf{X}_{t}})\,dW^{Y,i}_{t}\\ d\overline{X}^{i}_{t}\ &=\ b(t,\overline{X}^{i}_{t},\overline{\mu}_{t})\,dt\ +\ \sigma(t,\overline{X}^{i}_{t},\overline{\mu}_{t})\,dW^{X,i}_{t},\\ d\overline{Y}^{i}_{t}\ &=\ b(t,\overline{Y}^{i}_{t},\overline{\rho}_{t})\,dt\ +\ \sigma(t,\overline{Y}^{i}_{t},\overline{\rho}_{t})\,dW^{Y,i}_{t},\end{split}$ (87)
for $i\in\{1,\ldots,N\}$, with identical initial conditions $X^{i}_{0}=Y^{i}_{0}=\overline{X}^{i}_{0}=\overline{Y}^{i}_{0}=X_{0}$, we have
$\lim_{N\to\infty}\sup_{1\leq i\leq N}\mathbb{E}\left(\sup_{0\leq t\leq T}\lvert X^{i}_{s}-\overline{X}^{i}_{s}\rvert^{2}+\sup_{0\leq t\leq T}\lvert Y^{i}_{s}-\overline{Y}^{i}_{s}\rvert^{2}\right)\ =\ 0,$ (88)
@@
$\sup_{1\leq i\leq N}\left(\mathbb{E}\sup_{0\leq t\leq T}\lvert X^{i}_{t}-\overline{X}^{i}_{t}\rvert^{2}\ +\ \mathbb{E}\sup_{0\leq t\leq T}\lvert Y^{i}_{t}-\overline{Y}^{i}_{t}\rvert^{2}\right)\ \leq\ \frac{c}{N},$ (89)
@@
$\begin{split}&\leq\ c\int^{t}_{0}\mathbb{E}\lvert b(s,X^{i}_{s},\delta_{\mathbf{Y}_{s}})-b(s,\overline{X}^{i}_{s},\delta_{\overline{\mathbf{X}}_{s}})\rvert^{2}\,ds\\ &\qquad\qquad+\ c\int^{t}_{0}\mathbb{E}\lvert b(s,\overline{X}^{i}_{s},\delta_{\overline{\mathbf{X}}_{s}})-b(s,\overline{X}^{i}_{s},\overline{\mu}_{s})\rvert^{2}\,ds\\ &+\ c\int^{t}_{0}\mathbb{E}\lvert\sigma(s,X^{i}_{s},\delta_{\mathbf{Y}_{s}})-\sigma(s,\overline{X}^{i}_{s},\delta_{\overline{\mathbf{X}}_{s}})\rvert^{2}\,ds\\ &\qquad\qquad+\ c\int^{t}_{0}\mathbb{E}\lvert\sigma(s,\overline{X}^{i}_{s},\delta_{\overline{\mathbf{X}}_{s}})-\sigma(s,\overline{X}^{i}_{s},\overline{\mu}_{s})\rvert^{2}\,ds,\end{split}$ (92)
with an similar inequality holding for $\mathbb{E}\sup_{0\leq s\leq t}\lvert Y^{j}_{s}-\overline{Y}^{j}_{s}\rvert^{2}$.
Using the triangle inequality, and using a bound on the Wasserstein distance between two Dirac measures (Equation 1.24 [7])
$\mathcal{W}_{2}^{2}(\delta_{\mathbf{Y}_{s}},\delta_{\overline{\mathbf{X}}_{s}})\ \leq\ \mathcal{W}_{2}^{2}(\delta_{\mathbf{Y}_{s}},\delta_{\overline{\mathbf{Y}}_{s}})+\mathcal{W}_{2}^{2}(\delta_{\overline{\mathbf{Y}}_{s}},\delta_{\overline{\mathbf{X}}_{s}})\leq\ \frac{1}{N}\sum_{i=1}^{N}\lvert Y_{s}^{i}-\overline{Y}_{s}^{i}\rvert^{2}+\mathcal{W}_{2}^{2}(\delta_{\overline{\mathbf{Y}}_{s}},\delta_{\overline{\mathbf{X}}_{s}}).$ (93)
Since the random processes $\overline{\mathbf{Y}}_{t}$ and $\overline{\mathbf{X}}_{t}$ start with the same initial condition, existence and uniqueness of strong solutions to the McKean–Vlasov process implies uniquness in law, $\mathcal{W}_{2}^{2}(\delta_{\overline{\mathbf{Y}}_{s}},\delta_{\overline{\mathbf{X}}_{s}})=0$. Thus, by the exchangeability of the couples $(Y^{i},\overline{Y}^{i})$, the Lipschitz assumption of $b$ and $\sigma$ bounds the first and third term of (LABEL:appendix:reandom_eqquation) by
$c\int^{t}_{0}\left(\mathbb{E}\sup_{0\leq r\leq s}\lvert X_{r}^{i}-\overline{X}^{i}_{r}\rvert^{2}\ +\ \mathbb{E}\sup_{0\leq r\leq s}\lvert Y^{i}_{r}-\overline{Y}^{i}_{r}\rvert^{2}\right)\,ds.$ (94)
Combining the inequality for $\mathbb{E}\sup_{0\leq s\leq t}\lvert Y^{j}_{s}-\overline{Y}^{j}_{s}\rvert^{2}$, Grönwall’s lemma yields and the Lipschitz assumption yields
$\mathbb{E}\sup_{0\leq t\leq T}\lvert X^{i}_{t}-\overline{X}^{i}_{t}\rvert^{2}\ +\ \mathbb{E}\sup_{0\leq t\leq T}\lvert Y^{j}_{t}-\overline{Y}^{j}_{t}\rvert^{2}\$$\leq\ c\int^{T}_{0}\mathbb{E}\lvert b(s,\overline{X}^{i}_{s},\delta_{\overline{\mathbf{X}}_{s}})-b(s,\overline{X}^{i}_{s},\overline{\mu}_{s})\rvert^{2}\,ds$ (95)
@@
$\quad+\ c\int^{T}_{0}\mathbb{E}\lvert b(s,\overline{Y}^{i}_{s},\delta_{\overline{\mathbf{Y}}_{s}})-b(s,\overline{Y}^{i}_{s},\overline{\rho}_{s})\rvert^{2}\,ds$ (97)
$\qquad\qquad+\ c\int^{T}_{0}\mathbb{E}\lvert\sigma(s,\overline{Y}^{i}_{s},\delta_{\overline{\mathbf{Y}}_{s}})-\sigma(s,\overline{Y}^{i}_{s},\overline{\rho}_{s})\rvert^{2}\,ds$ (98)
$\leq c\int^{T}_{0}\left(\mathbb{E}\mathcal{W}_{2}^{2}(\delta_{\overline{\mathbf{X}}_{s}},\overline{\mu}_{s})\ +\ \mathbb{E}\mathcal{W}_{2}^{2}(\delta_{\overline{\mathbf{Y}}_{s}},\overline{\rho}_{s})\right)\,ds.$ (99)
Since $\overline{X}_{s}^{i},\overline{Y}_{s}^{j}$ are all independent random variables with common law $\overline{\mu}_{s},\overline{\rho}_{s}$, it follows by Lemma 2 and Lebesgue’s dominated convergence theorem that
$\lim_{N\to\infty}\sup_{1\leq i\leq N}\mathbb{E}\left(\sup_{0\leq t\leq T}\lvert X^{i}_{t}-\overline{X}^{i}_{t}\rvert^{2}+\sup_{0\leq t\leq T}\lvert Y^{i}_{t}-\overline{Y}^{i}_{t}\rvert^{2}\right)\ =\ 0.$ (100)
@@
$\qquad\qquad\qquad\qquad\cdot\bigg[\int b(s,\overline{X}^{i}_{s},y)\overline{\mu}_{s}(dy)-b(s,\overline{X}^{i},\delta_{\overline{X}^{i}_{s}})\bigg]\bigg).$ (103)
@@
$\sup_{1\leq i\leq N}\left(\mathbb{E}\sup_{0\leq t\leq T}\lvert X^{i}_{t}-\overline{X}^{i}_{t}\rvert^{2}\ +\ \mathbb{E}\sup_{0\leq t\leq T}\lvert Y^{i}_{t}-\overline{Y}^{i}_{t}\rvert^{2}\right)\ \leq\ \frac{c}{N}.$ (104)
@@ A.5 Adaptive MALA
The Metropolis--Hastings acceptance ratio is chosen as